OpenStudy (anonymous):

Trigonometry : Eliminate x : a cosec x+ b sec x + c =0, p cosec x +q sec x +r =0

OpenStudy (tkhunny):

Eliminate y \(\dfrac{a}{y} + by + c = 0\) \(\dfrac{p}{y} + qy + r = 0\)

OpenStudy (anonymous):

pls show me how to do @tkhunny if possible !!

OpenStudy (anonymous):

\[\frac{ \csc x }{ br-cq }=\frac{ \sec x }{ cp-ar }=\frac{ 1 }{ aq-bp }\] \[\csc x=\frac{ br-cq }{aq-bp },\sec x=\frac{ cp-ar }{ aq-bp }\] \[\sin x=\frac{ aq-bp }{ br-cq },\cos x=\frac{ aq-bp }{ cp-ar }\] square and add ,you get the result.

OpenStudy (anonymous):

@surjithayer ... thanks :)

OpenStudy (anonymous):

YW